# Deep Metric Imitation Learning for Stable Motion Primitives

### v2

@@ 3.5 Stability Analysis through Comparison Functions
We employ comparison functions, namely class-$\mathcal{KL}$ functions, to prove that the proposed method, based on the triplet loss, enforces global asymptotic stability at the equilibrium $x_{\mathrm{g}}$. These functions are used to formulate a general approach for stability analysis in the sense of Lyapunov. As described in [41, 42], these functions are defined as follows:
@@
Definition 2 ($-\mathcal{L}$ [^2] Please note that there is a distinction in the use of the symbol $\mathcal{L}$. While it represents the $-\mathcal{L}$ functions, it is also used to denote the set $\mathcal{L}$ introduced in Sec. 3.3. In subsequent sections of the paper, the $\mathcal{L}$ referring to $-\mathcal{L}$ functions is exclusively used in the form $\mathcal{KL}$. function).
A continuous function ${\sigma:\mathbb{R}_{\geq 0}\to\mathbb{R}_{>0}}$, is said to belong to class $\mathcal{L}$ if it is weakly decreasing [^3] We use this term to denote functions that either remain constant or strictly decrease within any interval of their domain. While these functions are often referred to as decreasing or non-increasing, we find these terms potentially ambiguous. and $\lim_{s\to\infty}\sigma(s)=0$.
@@
A function $\beta:[0,a)\times\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$, for $a\in\mathbb{R}_{>0}$, is said to belong to class $\mathcal{KL}$ if:
@@ 4 Methodology
We aim to learn motions from demonstrations, which are modeled as nonlinear time-invariant dynamical systems. In this context, we introduce the Policy via neUral Metric leArning (PUMA) framework, which learns the parametrized function $f^{\mathcal{T}}_{\theta}$ from human demonstrations. Furthermore, it guarantees that the goal of the motion $x_{\mathrm{g}}$ is a globally asymptotically stable equilibrium of this dynamical system, while maintaining accuracy with respect to the demonstrations. To achieve this, we augment the Imitation Learning (IL) problem with the loss $\ell_{\text{stable}}$ designed to enforce stability in the learned dynamical system based on the triplet loss. Hence, our framework minimizes:
$\ell_{\text{TIL}}=\ell_{\text{IL}}+\lambda\ell_{\text{stable}},$ (8)
where $\lambda\in\mathbb{R}_{>0}$ is a weight factor. We denote $\ell_{\text{TIL}}$ as the Triplet Imitation Learning (TIL) loss.
@@ 4.1 Behavioral Cloning
Figure 3: Illustration of the behavioral cloning loss computation. Starting from an initial condition $x_{0}$, the system $f^{\mathcal{T}}_{\theta}$ evolves to various time instants via $\Phi^{x}_{\theta}$. At each instant, the estimated state is compared with a demonstrated state. The red arrows show the path used to update the DNN weights using BPTT, for each $x_{t}$ as a function of $x_{0}$.
To minimize $\ell_{\text{IL}}$, and thereby address (3a), we adopt the behavioral cloning loss used in [9]. This loss enables the optimization of a deterministic value by representing the dynamical system as the mean of a Gaussian distribution with a fixed covariance. Furthermore, it mitigates covariate shift by minimizing the multi-step error over trajectory segments $\mathcal{H}^{i}$ using backpropagation through time (BPTT).
In every training iteration, we sample a batch $\mathcal{B}^{i}$ of trajectory segments $\mathcal{H}^{i}$ from the dataset $\mathcal{D}$. These segments can start at any point within a given demonstration, with the start time defined as $t=0$. Then, by introducing the evolution function $\Phi^{x}_{\theta}(t,x_{0}):\mathbb{R}_{\geq 0}\times\mathcal{T}\to\mathcal{T}$, which defines the estimated value of $x_{t}$ by integrating $f^{\mathcal{T}}_{\theta}$ between 00 and $t$, with initial condition $x_{0}$, we have
$\ell_{\text{IL}}=\sum_{\mathcal{H}^{i}\in\mathcal{B}^{i}}\sum_{(t,x^{*}_{t})\in\mathcal{H}^{i}}||x^{*}_{t}-\Phi^{x}_{\theta}(t,x_{0})||_{2}^{2},$ (9)
where the states $x^{*}_{t}$ along the trajectory segment $\mathcal{H}^{i}$ serve as labels for the states predicted by the DNN from the initial condition $x_{0}$ using $\Phi^{x}_{\theta}(t,x_{0})$. Note that the initial condition is obtained from $\mathcal{H}^{i}$, i.e., $x_{0}=x^{*}_{0}$. Fig. 3 illustrates the computation of this loss for one trajectory.
@@ 4.2.1 Reformulating the Stability Conditions
The stability conditions of Theo. 1 involve three dynamical systems: $f^{\mathcal{T}}_{\theta}$, $f^{\mathcal{T}\to\mathcal{L}}_{\theta}$, and $f^{\mathcal{L}}$. Note, however, that its first condition, namely $f_{\theta}^{\mathcal{T}\to\mathcal{L}}=f^{\mathcal{L}}$ $\forall x_{t}\in\mathcal{T}$, essentially states that $f_{\theta}^{\mathcal{T}\to\mathcal{L}}$ must exhibit global asymptotic stability. In other words, if the behavior of $f_{\theta}^{\mathcal{T}\to\mathcal{L}}$ is identical to that of another system, $f^{\mathcal{L}}$, for which stability is verified, then the stability of $f_{\theta}^{\mathcal{T}\to\mathcal{L}}$ is also verified. Nonetheless, if we can verify its stability through a different method, these conditions can be more generally written as follows:
@@ 4.2.2 Surrogate Stability Conditions
Theo. 3 introduces two conditions. If these are met within $f^{\mathcal{T}}_{\theta}$, the system will possess a globally asymptotically stable equilibrium at $x_{\mathrm{g}}$. However, these conditions do not offer a method for optimizing DNNs in a manner that ensures their fulfillment. Therefore, we introduce the surrogate stability conditions of Theo. 3. These conditions, when met, imply that the stability conditions of Theo. 3 are satisfied. Unlike the stability conditions, the surrogate conditions can be transformed into a loss function for optimizing the DNN, namely $\ell_{\text{stable}}$, which enforces their satisfaction.
To formulate the surrogate stability conditions, we note that Theo. 3 can be expressed in terms of relative distances. To facilitate this, we define the distance between any given latent state $y_{t}$ and a fixed target state $y_{\mathrm{g}}$ as $d_{t}=d(y_{\mathrm{g}},y_{t})={||y_{\mathrm{g}}-y_{t}||}$. Then, we formally introduce the conditions as:
@@
To prove this theorem, we demonstrate that if the surrogate stability conditions are satisfied, then the stability conditions from Theo. 3 must also hold. This leads to $x_{\mathrm{g}}$ being a globally asymptotically stable equilibrium of $f^{\mathcal{T}}_{\theta}$.
   a)  Second stability condition of Theo. 3
@@
Let us begin by analyzing the fulfillment of the second stability condition from Theo. 3. This condition states that only $x_{\mathrm{g}}$ can map to $y_{\mathrm{g}}$ via $\psi_{\theta}$. Then, since $y_{\mathrm{g}}$ is defined as $\psi_{\theta}(x_{\mathrm{g}})$, we need to establish that no other $x_{t}$ maps to $y_{\mathrm{g}}$. To achieve this, we must note that both $x_{t}$ and $x_{0}$ belong to the same state space $\mathcal{T}$. Therefore, showing that this statement holds $\forall x_{0}\in\mathcal{T}$ implies that it also holds $\forall x_{t}\in\mathcal{T}$.
@@
Now, assume, for the sake of contradiction, that there exists some $x_{0}$ such that $x_{0}\neq x_{\mathrm{g}}$ and $y_{0}=\psi_{\theta}(x_{0})=y_{\mathrm{g}}$. Also assume that the surrogate conditions 1 and 2 hold. Then, according to condition 2, $y_{t}\neq y_{t+\Delta t}$, which contradicts condition 1. Consequently, the second stability condition must be satisfied.
   b)  First stability condition Theo. 3
@@
Let us now study the first stability condition. Our goal is to prove that $y_{\mathrm{g}}$ is a globally asymptotically stable equilibrium of $f_{\theta}^{\mathcal{T}\to\mathcal{L}}$ within the region $\mathcal{L}$. Surrogate condition 2 hints that the system’s stability might be verified using a Lyapunov candidate defined as a function of $d_{t}$. This is because the condition enforces the distance $d_{t}$ to strictly decrease within the interval defined by $\Delta t$. However, this does not necessarily imply that the Lyapunov candidate consistently strictly decreases with time, as it is possible for it to strictly increase locally while adhering to this condition, as illustrated in Fig. 4. Therefore, we propose to demonstrate the global asymptotic stability of $y_{\mathrm{g}}$ using Theo. 2. Specifically, we aim to do so by employing a class-$\mathcal{KL}$ upper bound $\beta$ that fulfills (7).
@@
Figure 4: Time evolution of two functions $\delta$, $\delta_{1}$ and $\delta_{2}$, starting with different initial conditions $y_{0}$, but same $d_{0}$. Both satisfy the surrogate stability conditions for $\Delta t=2$. Additonally, the values of $\delta^{\text{max}}$ and $\beta$, computed using these functions, are shown. In this representation, ${\delta=e^{-a\cdot t}\left(\sin^{2}\left(\omega t\right)+d_{0}\cdot\cos^{2}\left(\omega t\right)\right)}$ with $a=0.75$, $d_{0}=0.2$, and $\omega=[\pi,\pi/2]$.
@@
To achieve this, we first define an evolution function for the distance $d_{t}$, for a given $y_{0}$ and $t$, as $\delta:\mathcal{L}\times\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$, with $\delta(y_{0},t)=||y_{\mathrm{g}}-\Phi^{y}_{\theta}(t,y_{0})||$. Here, $\Phi^{y}_{\theta}$ represents the evolution function of $y_{t}$ under the dynamical system $f^{\mathcal{T}\to\mathcal{L}}_{\theta}$. Then, we can express the upper bound condition of $\beta$ as
@@
It is important to note that verifying all these properties for a given $\beta$ can lead to a lengthy proof. Thus, while we introduce $\beta$ and discuss the key concepts behind its design here, the complete proof and details are provided in Prop. 1, Appendix A.
@@
  1. i)
    Upper bound in time: First, we need to identify an upper bound that weakly decreases with time. For this purpose, we introduce a function that computes the maximum of $\delta$ over the window $[t,t+\Delta t]$ for any given $y_{0}$ and $t$, i.e.,
@@
    Clearly, this function serves as an upper bound for $\delta$. Moreover, this function must weakly decrease with time. Considering surrogate condition 2, which indicates that $\forall s\in[t,t+\Delta t]$, we have $\delta(y_{0},s+\Delta t)<\delta(y_{0},s)$, it follows that there is no $\delta$ greater than $\delta^{\text{max}}_{t+\Delta t}$ in the interval $[t+\Delta t,t+2\Delta t]$. By recursively extending this observation for every interval $[t+n\cdot\Delta t,t+(n+1)\cdot\Delta t]$, with $n\in\mathbb{N}$, we can conclude that $\delta^{\text{max}}_{t+\Delta t}$ weakly decreases with time.
  2. ii)
    Upper bound in space: The function $\delta^{\text{max}}_{t+\Delta t}(y_{0},t)$ provides an upper bound of $\delta$ for a given $y_{0}$. However, $\beta(d_{0},t)$ depends on $d_{0}$ rather than directly on $y_{0}$. To address this, for any specified $d_{0}$, we must ensure that $\beta\geq\delta$ for every $y_{0}$ located at this particular distance from the equilibrium. The set of $y_{0}$ fulfilling this condition can be defined as $\mathcal{Y}_{0}(d_{0})=\{y_{0}\in\mathcal{L}:||y_{\mathrm{g}}-y_{0}||=d_{0}\}$. Taking this into account, we can then introduce the following upper bound that depends on $d_{0}$:
@@
  3. iii)
    Strictly increasing/decreasing function: Similarly to $\delta^{\text{max}}_{t+\Delta t}$, the function $\delta^{\text{max}}$ also weakly decreases as a function of time (see Appendix A for details). This is not inherently problematic, as it verifies the properties of class-$\mathcal{KL}$ functions. However, $\beta$ must strictly increase as a function of $d_{0}$, and, for this to be the case in our formulation, $\beta$ must also strictly decrease as a function of time, $\forall d_{0}\in\mathcal{L}\setminus\{y_{\mathrm{g}}\}$.
    To achieve this, we treat $\beta$ as the evolution function of a first-order linear dynamical system with state $z_{t}$, using $\delta^{\text{max}}$ as the reference. This leads to the equation
@@
    where $\alpha<0$ and an initial condition $z_{0}$ greater than $\delta^{\text{max}}_{0}=\delta^{\text{max}}(d_{0},0)$ ensures that $\beta$ strictly decreases with time and remains above $\delta^{\text{max}}$ (and, consequently, above $\delta$) for all $d_{0}\in\mathcal{L}\setminus\{y_{\mathrm{g}}\}$. By choosing $z_{0}=\delta^{\text{max}}_{0}+d_{0}$, $\beta$ can then be expressed as
@@
Based on Prop. 1, Appendix A, we can confirm that our formulation for $\beta$ is a valid class-$\mathcal{KL}$ upper bound of $\delta$. This indicates that the second stability condition of Theo. 3 is satisfied, concluding our proof. ∎∎

### v3

@@ 3.5 Stability Analysis through Comparison Functions
In this work, we employ comparison functions, namely class-$\mathcal{KL}$ functions, to prove global asymptotic stability at the equilibrium $x_{\mathrm{g}}$. These functions are used to formulate a general approach for stability analysis in the sense of Lyapunov. As described in [43, 44], these functions are defined as follows:
@@
Definition 2 ($-\mathcal{L}$ function [^2] Note that the symbol $\mathcal{L}$ is used in two distinct contexts. While it represents the $-\mathcal{L}$ functions, it is also used to denote the set $\mathcal{L}$ introduced in Sec. 3.3. In subsequent sections of the paper, the $\mathcal{L}$ referring to $-\mathcal{L}$ functions is exclusively used in the form $\mathcal{KL}$.).
A continuous function ${\sigma:\mathbb{R}_{\geq 0}\to\mathbb{R}_{>0}}$, is said to belong to class $\mathcal{L}$ if it is weakly decreasing [^3] We use this term to denote functions that either remain constant or strictly decrease within any interval of their domain. and $\lim_{s\to\infty}\sigma(s)=0$.
@@
A function $\beta:[0,a)\times\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$, for $a\in\mathbb{R}_{>0}$, is said to belong to class-$\mathcal{KL}$ if:
@@ 4 Methodology
We introduce the Policy via neUral Metric leArning (PUMA) framework, which learns motion primitives from human demonstrations parametrized as the dynamical system $f^{\mathcal{T}}_{\theta}$. Furthermore, it enforces the goal of the motion $x_{\mathrm{g}}$ to be a globally asymptotically stable equilibrium of $f^{\mathcal{T}}_{\theta}$ while maintaining accuracy with respect to the demonstrations. To achieve this, we augment the Imitation Learning (IL) problem with the stability-enforcing loss $\ell_{\text{stable}}$. Hence, our framework minimizes the loss:
$\ell_{\scriptscriptstyle\mathrm{PUMA}}=\ell_{\text{IL}}+\lambda\ell_{\text{stable}},$ (8)
where $\ell_{\text{IL}}$ is an imitation learning loss and $\lambda\in\mathbb{R}_{>0}$ is a weight factor.
@@ 4.1 Behavioral Cloning
Figure 3: Illustration of the behavioral cloning loss computation. Starting from an initial condition $x_{0}$, the system $f^{\mathcal{T}}_{\theta}$ evolves to various time instants via $\Phi^{x}_{\theta}$. At each instant, the estimated state is compared with a demonstrated state. The red arrows show the gradient path used to update the DNN’s weights using BPTT.
To minimize $\ell_{\text{IL}}$, and thereby address (3a), we adopt the behavioral cloning loss used in [9]. This loss tackles (3a) by modeling the output of the deterministic dynamical system $f^{\mathcal{T}}_{\theta}$ as the mean of a Gaussian distribution with fixed covariance. Furthermore, it mitigates covariate shift by minimizing the multi-step error over trajectory segments using backpropagation through time (BPTT), as depicted in Fig. 3.
In every training iteration, we sample a batch $\mathcal{B}^{i}$ of trajectory segments $\mathcal{H}^{i}$ from the dataset $\mathcal{D}$. These segments can start at any point within a given demonstration, with the start time defined as $t=0$. Then, by introducing the evolution function $\Phi^{x}_{\theta}(t,x_{0}):\mathbb{R}_{\geq 0}\times\mathcal{T}\to\mathcal{T}$, which defines the value of $x_{t}$ by integrating $f^{\mathcal{T}}_{\theta}$ between 00 and $t$, with initial condition $x_{0}$, we can construct the loss
$\ell_{\text{IL}}=\sum_{\mathcal{H}^{i}\in\mathcal{B}^{i}}\sum_{(t,x^{*}_{t})\in\mathcal{H}^{i}}||x^{*}_{t}-\Phi^{x}_{\theta}(t,x_{0})||_{2}^{2}.$ (9)
Here, the states $x^{*}_{t}$ along the trajectory segment $\mathcal{H}^{i}$ serve as labels for the states predicted by the DNN from the initial condition $x_{0}$ using $\Phi^{x}_{\theta}(t,x_{0})$. Note that the initial condition is obtained from $\mathcal{H}^{i}$, i.e., $x_{0}=x^{*}_{0}$.
@@ 4.2.1 Reformulating the Stability Conditions
The stability conditions of Theorem 1 involve three dynamical systems: $f^{\mathcal{T}}_{\theta}$, $f^{\mathcal{T}\to\mathcal{L}}_{\theta}$, and $f^{\mathcal{L}}$. Note, however, that its first condition, namely, $f_{\theta}^{\mathcal{T}\to\mathcal{L}}=f^{\mathcal{L}}$ $\forall x_{t}\in\mathcal{T}$, essentially states that $f_{\theta}^{\mathcal{T}\to\mathcal{L}}$ must exhibit global asymptotic stability. In other words, if the behavior of $f_{\theta}^{\mathcal{T}\to\mathcal{L}}$ is identical to that of another system, $f^{\mathcal{L}}$, for which stability is verified, then the stability of $f_{\theta}^{\mathcal{T}\to\mathcal{L}}$ is also verified. Nonetheless, if we can enforce its stability through a different method, these conditions can be more generally written as follows:
@@ 4.2.2 Surrogate Stability Conditions
Theorem 3 introduces stability conditions for $f^{\mathcal{T}}_{\theta}$; nevertheless, it does not specify how to enforce these conditions in the system. Therefore, we introduce the surrogate stability conditions of Theorem 3. These conditions, when met, imply that the stability conditions of Theorem 3 are also satisfied. Unlike the stability conditions, the surrogate conditions can be directly transformed into a specific loss function, $\ell_{\text{stable}}$, for optimizing the DNN to enforce their satisfaction.
To formulate the surrogate stability conditions, we note that Theorem 3 can be expressed in terms of relative distances. We define the distance between any given latent state $y_{t}$ and the goal state $y_{\mathrm{g}}$ as $d_{t}=d(y_{\mathrm{g}},y_{t})=\|y_{\mathrm{g}}-y_{t}\|$. Then, according to Condition 1 of Theorem 3, which addresses global asymptotic stability, the value of $d_{t}$ should, generally speaking, decrease over time. Moreover, the second condition specifies that the value of $d_{t}$ should remain constant only for $y_{\mathrm{g}}=\psi_{\theta}(x_{\mathrm{g}})$. Formally, we introduce the conditions as:
@@
To prove this theorem, we demonstrate that if the surrogate stability conditions are satisfied, then the stability conditions from Theorem 3 must also hold. This leads to $x_{\mathrm{g}}$ being a globally asymptotically stable equilibrium of $f^{\mathcal{T}}_{\theta}$.
   a)  Second stability condition of Theorem 3
@@
Let us begin by analyzing the fulfillment of the second stability condition from Theorem 3. This condition states that only $x_{\mathrm{g}}$ can map to $y_{\mathrm{g}}$ via $\psi_{\theta}$. Then, since $y_{\mathrm{g}}$ is defined as $\psi_{\theta}(x_{\mathrm{g}})$, we need to establish that no other $x_{t}$ maps to $y_{\mathrm{g}}$. To achieve this, we note that both $x_{t}$ and $x_{0}$ belong to the same state space $\mathcal{T}$. Therefore, showing that this statement holds $\forall x_{0}\in\mathcal{T}$ implies that it also holds $\forall x_{t}\in\mathcal{T}$.
@@
Now, suppose for a contradiction that there exists some $x_{0}$ such that $x_{0}\neq x_{\mathrm{g}}$ and $y_{0}=\psi_{\theta}(x_{0})=y_{\mathrm{g}}$. Then, according to the second surrogate condition, $d_{t}>d_{t+\Delta t}$, which contradicts the first surrogate condition. Consequently, the second stability condition must be satisfied.
   b)  First stability condition Theorem 3
@@
Let us now study the first stability condition. Our goal is to prove that $y_{\mathrm{g}}$ is a globally asymptotically stable equilibrium of $f_{\theta}^{\mathcal{T}\to\mathcal{L}}$ within the region $\mathcal{L}$, where the system is defined. Surrogate condition 2 hints that the system’s stability could be verified through a Lyapunov candidate defined using $d_{t}$. This is because the condition enforces the distance $d_{t}$ to strictly decrease within the interval defined by $\Delta t$. However, this does not necessarily imply that the Lyapunov candidate strictly decreases with time, as it is possible for it to strictly increase locally while adhering to this condition, as exemplified in Fig. 4. As a result, we proceed to demonstrate the global asymptotic stability of $y_{\mathrm{g}}$ using Theorem 2. Specifically, we aim to do so by employing a class-$\mathcal{KL}$ upper bound $\beta$ that fulfills (7).
@@
Figure 4: Time evolution of two functions $\delta$, $\delta_{1}$ and $\delta_{2}$, starting with different initial conditions $y_{0}$, but same $d_{0}$. Both satisfy the surrogate stability conditions with ${\Delta t=2}$. Additionally, the values of $\delta^{\text{max}}$ and $\beta$, computed using these functions, are shown. In this representation, ${\delta=e^{-a\cdot t}\left(\sin^{2}\left(\omega t\right)+d_{0}\cdot\cos^{2}\left(\omega t\right)\right)}$ with $a=0.75$, $d_{0}=0.2$, and $\omega=[\pi,\pi/2]$.
@@
To achieve this, we first define an evolution function for the distance $d_{t}$, for a given $y_{0}$ and $t$, as $\delta:\mathcal{L}\times\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$, with $\delta(y_{0},t)=||y_{\mathrm{g}}-\Phi^{y}_{\theta}(t,y_{0})||$. Here, $\Phi^{y}_{\theta}$ represents the evolution function of $y_{t}$ under the dynamical system $f^{\mathcal{T}\to\mathcal{L}}_{\theta}$. Then, we can express the upper bound $\beta$ as
@@
It is important to note that verifying all these properties can lead to a lengthy proof. Thus, while we introduce $\beta$ and discuss the key concepts behind its design here, the complete proof and details are provided in Proposition 1, Appendix A.
@@
  1. 1.
    Upper bound in time: First, we need to identify a $\beta$ that weakly decreases with time. For this purpose, we introduce a function that computes the maximum of $\delta$ over the window $[t,t+\Delta t]$ for any given $y_{0}$ and $t$, i.e.,
@@
    Clearly, this function serves as an upper bound for $\delta$. Moreover, this function must weakly decrease with time. Considering surrogate condition 2, which indicates that $\forall s\in[t,t+\Delta t]$, we have $\delta(y_{0},s+\Delta t)<\delta(y_{0},s)$, it follows that there is no $\delta$ greater than $\delta^{\text{max}}_{t+\Delta t}$ in the interval $[t+\Delta t,t+2\Delta t]$. By extending this observation for every interval $[t+n\cdot\Delta t,t+(n+1)\cdot\Delta t]$, with $n\in\mathbb{N}$, we can conclude that $\delta^{\text{max}}_{t+\Delta t}$ weakly decreases with time.
  2. 2.
    Upper bound in space: The function $\delta^{\text{max}}_{t+\Delta t}(y_{0},t)$ provides an upper bound of $\delta$ for a given $y_{0}$. However, $\beta(d_{0},t)$ depends on $d_{0}$ rather than directly on $y_{0}$. To address this, for any specified $d_{0}$, we must ensure that $\beta\geq\delta$ for every $y_{0}$ located at this particular distance from the equilibrium. The set of initial conditions $y_{0}$ fulfilling this condition can be defined as $\mathcal{Y}_{0}(d_{0})=\{y_{0}\in\mathcal{L}:||y_{\mathrm{g}}-y_{0}||=d_{0}\}$. Considering this, we can introduce an upper bound dependent on $d_{0}$ by computing the maximum of each $\delta^{\text{max}}_{t+\Delta t}(y_{0},t)$, where $y_{0}\in\mathcal{Y}_{0}$:
@@
  3. 3.
    Strictly increasing/decreasing function: Similarly to $\delta^{\text{max}}_{t+\Delta t}$, the function $\delta^{\text{max}}$ also weakly decreases as a function of time (see Appendix A for details). This is not inherently problematic, as it verifies the properties of class-$\mathcal{KL}$ functions. However, $\beta$ must strictly increase as a function of $d_{0}$, and, for this to be the case, in our formulation, $\beta$ must also strictly decrease as a function of time, $\forall d_{0}\in\mathcal{L}\setminus\{y_{\mathrm{g}}\}$.
    To achieve this, we consider $\beta$ as the evolution function of a first-order linear dynamical system with state $z_{t}$, using $\delta^{\text{max}}$ as the reference. This leads to the equation
@@
    where $\alpha<0$. With an initial condition $z_{0}$ greater than $\delta^{\text{max}}_{0}=\delta^{\text{max}}(d_{0},0)$, we ensure that $\beta$ strictly decreases with time. Moreover, $\beta$ remains above $\delta^{\text{max}}$, and consequently above $\delta$, for all $d_{0}\in\mathcal{L}\setminus\{y_{\mathrm{g}}\}$. Hence, by choosing ${z_{0}=\delta^{\text{max}}_{0}+d_{0}}$, we ensure that $\beta$ exhibits the desired increasing/decreasing properties. We can express $\beta$ as
@@
Based on Proposition 1, Appendix A, we can confirm that our formulation for $\beta$ is a valid class-$\mathcal{KL}$ upper bound of $\delta$. This indicates that the first stability condition of Theorem 3 is satisfied, concluding our proof. ∎